INAF queue observing, 2016 December 2 - 3

Observer: R. Carini, A. Marchetti
Telescope Operator: G. Bechetti
Support Astronomer: D. Thompson
Instrument: LUCI1, MODS1, MODS2

Summary

At the sunset the sky is partly cloudy, but storm clouds are fastly approaching. Snow is forecast at 60%. The TO does not open the enclosure.

LUCI2 is not available

01:24 It starts snowing

Night duration (12\xB0 el twilight) 18:10-6:00 ( 01:10-13.0 UT) = 11h 50m

Observing time: 0h0m (0%)

Weather time loss: 11 h 50m (100%)

technical time loss 0m (0%)

Data Summary

Project_RUN Objects Filter ExpTime Notes
DARKS LUCI1 IOVINO, PICONCELLI, RICCI, MOSCADELLI      
FLATS LUCI1 IOVINO, PICONCELLI, RICCI, MOSCADELLI      
ARCS LUCI1 IOVINO, PICONCELLI, RICCI, MOSCADELLI      
FLATS (LS) MODS1+MODS2 GRAZIAN, PANCINO      
ARC (LS) MODS1+MODS2 GRAZIAN, PANCINO      
FLATS STANDARD MODS1+MODS2 GRAZIAN, PANCINO      
BIAS 8k MODS1+MODS2        
BIAS 3k MODS1+MODS2        

INAF queue observing, 2016 December 3 - 4

Observer: R. Carini, A. Marchetti
Telescope Operator: G. Bechetti
Support Astronomer: D. Thompson
Instrument: LUCI1, MODS1, MODS2

Summary

At sunset the sky is clear, but humidity is at 95%, and temperature very near to dew point.

LUCI2 is not available

00:55 humidity 100%, temperature at the dew point.
the dome is closed
during the whole night the seeing was very bad, oscillating between 1.5'' e 2.5'', with peaks to 3''

Night duration (12\xB0 el twilight) 18:10-6:00 ( 01:10-13.0 UT) = 11h 50m

Observing time: 6h 15m (53%)

Weather time loss: 5h 35m (47%)

technical time loss 0m (0%)

Data Summary

Project_RUN Objects Filter ExpTime Notes
LUCI 1 Antoniucci, Marziani B DARKS      
LUCI 1 Antoniucci, Marziani B FLATS      
LUCI 1 Antoniucci, Marziani B ARCS      
MODS 1-2 Pancino Pandas-17     seeing out of spec, misunderstanding between observers, the target has been already taken
MODS 1-2 Antoniucci V1118Ori dual grating 20 m spec COMPLETED
MODS 1-2 Antoniucci V1118Ori blue : g_sdss; red: r_sdss 2m imaging COMPLETED
MODS 1-2 Standard Feige34 dual grating    
LUCI1 Marziani J SDSS J1054+2536 G210HiRes H 0.5 h seeing out of spec, the frames taken in twilight from the third
LUCI1 SKy FLAT   Ks, H    

INAF queue observing, 2016 December 4 - 5

Observer: R. Carini, A. Marchetti
Telescope Operator: G. Bechetti
Support Astronomer: O. Kuhn on call
Instrument: LUCI1, LBC

Summary

LUCI2 is not available

At sunset the sky is clear. There is wind coming from the north-east. We need to choose targets towards the south.

Night duration (12\xB0 el twilight) 18:10-6:00 ( 01:10-13.00 UT) = 11h 50m

Observing time: 9h 30m (80%)

Weather time loss: 0h 0m (0%)

technical time loss 2h 20m (20%)

Data Summary

Project_RUN | Objects | Filter | ExpTime | Notes |
Marziani B SDSS J0008-1039 H 18m seeing very variable. taken the last two images to complete the target, plus one due to bad seeing.
telluric Marzini B HIP1378 H    
Antoniucci ZXtau zj 0.2 h miss HK filter and ima
telluric Iovino HIP11809 HKspec    
Iovino M004XXLrev HKspec 1 h last frame with 1.4 < airmass <1.43 and 0.85'' < seeing< 1.00''
telluric Antoniucci HIP30393 zj    
Castellano ob. 17 MACJ0717 Uspec, V COMPLETED
Castellano ob. 18 MACJ0717 Uspec, V COMPLETED
Castellano ob. 19 MACJ0717 Uspec, V COMPLETED, but seeing oscillating. CHECK
photometric star SA101 blue:Uspec, g, red:V,r    
Bellazzini I 1437+12.9 blue:g,red:r 0.2h COMPLETED
Bellazzini J 147.0+07.1 blue:g,red:r 0.2h COMPLETED
photometric star SA104 blue:Uspec, g, red:V,r
DAlessandro Pal14 1 image in blue, 3 in red, during twiligth
CALIBRATION sky flats for LBC blue U_spec, g, red V, r
CALIBRATION bias for LBC

INAF queue observing, 2016 December 5 - 6

Observer: R. Carini, A. Marchetti
Telescope Operator: G. Bechetti
Support Astronomer: O. Kuhn on call
Instrument: LUCI1, MODS1, MODS2

Summary

LUCI2 is not available

At sunset the sky is cloudy and there is west-southwest wind. For all the night there was clouds and strong wind until 3 a.m. (limiting the observation to targets in north-northeast direction), with horrible and strongly variable seeing, with peaks of 4". We experienced technical problems with the presets, both with LUCI and MODS: some scripts couldn't be sent, it is not clear if it is a telescope issue or a scripts issue, or something related to the LUCI2 to LUCI1 conversion.

Night duration (12\xB0 el twilight) 18:10-6:00 ( 01:45-12.45 UT) = 11h 50m

Observing time: 7h 56m (67%)

Weather time loss: 1h 37m (13%)

technical time loss 2h 17m (20%)

Data Summary

Project_RUN Objects Filter ExpTime Notes
Antoniucci PV Cep zj 300 sec seeing variable 1.7''-2.8''
Antoniucci V1118Ori zJ+HK   COMPLETED but seeing strongly variable
telluric Antoniucci zj+ Hk    
Annibali DD068 RunB dual grating 1 h 1 ob completed in MODS1 only.
Annibali Standard Feige34 dual grating    
CALIBRATIONS Annibali MODS 1 DD068      
CALIBRATIONS Antoniucci MODS1        

INAF queue observing, 2016 December 6 - 7

Observer: R. Carini, A. Marchetti
Telescope Operator: S. Allanson
Support Astronomer: B. Rothberg on call
Instrument: MODS1, MODS2, LUCI1

Summary

LUCI2 is not available

At sunset there is strong southwest wind and some clouds.

Night duration (12\xB0 el twilight) 18:10-6:00 ( 01:45-12.45 UT) = 11h 50m

Observing time: 5h 02m (43%)

Weather time loss: 6h 48m (57%)

technical time loss 0m (0%)

Data Summary

Project_RUN Objects Filter ExpTime Notes
Pancino H12 Dual 0.5 h COMPLETED MODS1+MODS2
Massaro 3FGL60258.2+3555 prism, Red 0.6 h (300sec x6) COMPLETED MODS1+MODS2
Pancino H18 Dual 0.5 h (250sec x6 in each filter) COMPLETED MODS1+MODS2
Pancino H3 Dual 36 m (1100sec x2 in each filter) filter MODS1+MODS2,
Calibrations Annibali MODS1 flats and arcs for DD068_RunB Dual    
Calibrations Antoniucci MODS2 flats, arcs, imaging flats Dual    
Calibrations Massaro MODS1-MODS2 bias 3kx3k, 4kx4k Red    
Calibrations Massaro MODS1-MODS2 flats, arcs 1", 0.3" Red    
Standard Pancino MODS1-MODS2 g191b2b Dual   seeing oscillating, some clouds
Annibali MODS1-MODS2 DD068_RunB Dual 1200sec x2 parallactic angle out of spec
Annibali MODS1-MODS2 Feige34 Dual   parallactic angle out of spec

INAF queue observing, 2016 December 7 - 8

Observer: R. Carini, A. Marchetti
Telescope Operator: S. Allanson
Support Astronomer: B. Rothberg on call
Instrument: LUCi1

Summary

LUCI2 is not available

At sunset the sky is partly cloudy

Night duration (12\xB0 el twilight) 18:10-6:00 ( 01:45-12.45 UT) = 11h 50m

Observing time: 3h 34m (30%)

Weather time loss: 8h 16m (70%)

technical time loss 0m (0%)

Data Summary

Project_RUN Objects Filter ExpTime Notes
CALIBRATIONS Massaro MODS1-MODS2 Flats 1" Red    
Antoniucci LUCi1 PVCep zj 1200sec COMPLETED
Antoniucci LUCI1 PVCep HK 240sec COMPLETED
Antoniucci LUCI1 imaging PVCep     COMPLETED
Antoniucci LUCI1 telluric HIP111257 zJ+HK   0024 is not good, 0025 needs to be checked
CALIBRATIONS Massaro MODS1-MODS2 Flats 1.2" Red    
Ricci A2 telluric HIP41798 J   Last image taken with flexure compensation off
Ricci A2 target NGC2712 J 300sx6  
Ricci A2 target NGC2712 J 300sx6  
Ricci A2 target NGC2712 J 300sx2 missing 4 images of 300 sec for target completion

-- %USERSIG{RobertaCarini - 2016-12-07}%

Comments

 
Topic revision: r4 - 19 Dec 2016, RobertaCarini
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback